6x − 3y = 18
6x + 3y = −12
Solve using substitution

Answer by flame8855(424)

6x − 3y = 18
6x + 3y = −12
2x-y=6
y=2x-6
6x+3y=-12
6x+3(2x-6)=-12
6x+6x-18=-12
12x=6
x=1/2
y=-5